A train 150 m in length is going northwards at a speed of 10 m/s. A parrot flies at a speed of 5 m/s southwards, parallel to the railway track. What is the time taken by the parrot to cross the train?                                                                                                               

Grade:11

1 Answers

Khimraj
5 years ago
use concept of relative velocity
relative velocity = 5 – (-10) = 15m/s
relative displacement = 150
then time t = 150/15 = 10 sec
Hope it clears.